A General Electric 7½ % bond closed at 98. What is the current yield?

Hint: Review the definition of current yield for bond.
Answer #1

Calculation of current yield:

Current yield= coupon payment/current price*100

Let the face value be 100

Coupon payment= 100*0.075= 7.50

Current yield= 7.5/98*100= 7.65%

Current yield= 7.65%